# Nonparametric Estimation of Matching Efficiency and Mismatch in Labor Markets via Public Employment Security Offices in Japan, 1972-2024

### v1

@@ 5.5 Mismatch in 2012-2024
Finally, I compute the estimated nonparametric mismatch index in 2012-2024. Panel (a) of Figure 11 illustrates the estimated nonparametric mismatch index across 47 prefectures from 2014 to 2024. The index, representing the deviation of the actual number of hires from a social planner’s optimal one, demonstrates significant seasonal fluctuations throughout the period. Mismatch across prefectures increases to around 0.3. Notably, the mismatch index shows a cyclical pattern with peaks and troughs, suggesting seasonal factors or regional economic cycles influencing labor market dynamics. The mismatch level across prefectures is larger than previous Cobb-Douglas specification results such as Shibata (2020) and Higashi and Sasaki (2023).
Panel (b) of Figure 11 shows the mismatch index across different job categories over a similar period. Unlike the prefecture-based mismatch, this index exhibits a more consistent upward trajectory, with less pronounced seasonal variations. The mismatch index across job categories increases to 0.6, which is significantly larger than Shibata (2020) and Higashi and Sasaki (2023), which indicates that the standard Cobb-Douglas specification underestimates the mismatch index due to the absence of time-varying matching efficiency and elasticity. The steady increase in the mismatch index across job categories points to a growing dissonance between the skills offered by job seekers and the skills demanded by employers. This trend may reflect evolving industry demands, technological advancements, and the lag in the workforce’s adaptation to new skills. The sharper rise in the mismatch index post-2020 further suggests that the pandemic may have exacerbated skill mismatches, possibly due to shifts in industry dynamics, such as the acceleration of digital transformation and the differing impacts on various sectors.
@@ 6 Conclusion
I investigate how matching efficiency in the labor market via Public Employment Security Offices in Japan for unemployed workers changed during the period 1972-2024. By applying a novel nonparametric identification approach proposed by Lange and Papageorgiou (2020) to monthly data, I find that matching efficiency (normalized to 1972) exhibits a declining trend with notable fluctuations, consistent with the downward trends in job and worker finding rates. Finally, I extend the mismatch index proposed by Şahin et al. (2014) to the nonparametric version and find that the mismatch across job categories is more severe than across prefectures and the original Cobb-Douglas mismatch index is biased.

### v2

@@ 6.5 Mismatch in 2012-2024
Finally, I compute the estimated nonparametric mismatch index. Panel (a) of Figure 11 illustrates the estimated nonparametric mismatch index across 47 prefectures from 2014 to 2024. The index, representing the deviation of the actual number of hires from a social planner’s optimal one, demonstrates significant seasonal fluctuations throughout the period. Mismatch across prefectures increases to around 0.3. Notably, the mismatch index shows a cyclical pattern with peaks and troughs, suggesting seasonal factors or regional economic cycles influencing labor market dynamics. The mismatch level across prefectures is larger than previous Cobb-Douglas specification results such as Shibata (2020) (0.05-0.1) and Higashi and Sasaki (2023) (0.15-0.25).
Panel (b) of Figure 11 shows the mismatch index across different occupations over a similar period. Unlike the prefecture-based mismatch, this index exhibits a more consistent upward trajectory, with less pronounced seasonal variations. The mismatch index across occupations increases to 0.6, which is significantly larger than Shibata (2020) and Higashi and Sasaki (2023), which indicates that the standard Cobb-Douglas specification underestimates the mismatch index due to the absence of time-varying matching efficiency and elasticity. The steady increase in the mismatch index across occupations points to a growing dissonance between the skills offered by job seekers and the skills demanded by employers. This trend may reflect evolving industry demands, technological advancements, and the lag in the workforce’s adaptation to new skills. The sharper rise in the mismatch index post-2020 further suggests that the pandemic may have exacerbated skill mismatches, possibly due to shifts in industry dynamics, such as the acceleration of digital transformation and the differing impacts on various sectors.
@@ 7 Conclusion
I investigate how matching efficiency in the labor market via Public Employment Security Offices in Japan for unemployed workers changed during the period 1972-2024. By applying a novel nonparametric identification approach proposed by Lange and Papageorgiou (2020) to monthly data, I find that matching efficiency (normalized to 1972) exhibits a declining trend with notable fluctuations, consistent with the downward trends in job and worker finding rates. Finally, I extend the mismatch index proposed by Şahin et al. (2014) to the nonparametric version and find that the mismatch across occupations is more severe than across prefectures and the original Cobb-Douglas mismatch index is biased.